An exponential sequence is such that the third term minus the first term is 48. The fourth term minus the second term is 144. Find the common ratio, the first term, the sixth term
Answer by richwmiller(17219)

Tn = t * r^(n - 1)
T3 = t * r^2
T1 = t
t3-t1=48
(t * r^2)-t=48
T4 = t * r^(3)
T2 = t * r^(1)
t4-t2=144
t * r^(3)- t * r^(1)=144
r = 3, t1 = 6
T6 = 6 * 3^(5)
T6 = 1458